### v2

# Fixed Points of the Josephus Function via Fractional Base Expansions

Abstract
In this paper, we investigate some interesting properties of the fixed point sequence of the Josephus function $J_{3}$. First, we establish a connection between this sequence and the Chinese Remainder Theorem. Next, we identify a clear numerical pattern for the digits of two consecutive fixed points when they are written in a non-standard fractional number system in base $3/2$. This result allows us to derive a recursive procedure for determining the digits of their base $3/2$ expansions.
1 Introduction
This work discusses some interesting properties of the Josephus function $J_{3}$, which arises from the classical Josephus problem, a well-known mathematical riddle originating from a narrative by the Jewish-Roman historian Flavius Josephus (see [4, Book 3, Chapter 8, Part 7]). The problem is described as follows: Given a group of $n$ people seated at a round table, numbered consecutively clockwise from $1$ to $n$, begin counting from position $1$, skip two people, and eliminate the third. Continue this process with the remaining participants, always starting with the next person, counting two, and removing the third, proceeding clockwise until only one person remains. This person is called the survivor.
A natural question is where one should initially sit to avoid elimination. Let $J_{3}(n)$ denote the position of the survivor. Computing $J_{3}(n)$ directly for large $n$ can be costly, so it is useful to identify features of $J_{3}$ that can be exploited; see [3, 2].
Next, we introduce the notion of extremal points and explain their role in the structure of $J_{3}$.
Definition (High extremal points).
A high extremal point is an integer $n_{e}$ such that
$J_{3}(n_{e})\in\{n_{e}-1,n_{e}\}.$
High extremal points come in two types:
  1. (a)
    If $J_{3}(n_{e})=n_{e}$, then $n_{e}$ is a fixed point of $J_{3}$, denoted by $n_{p}$.
  2. (b)
    If $J_{3}(n_{e})=n_{e}-1$, then $n_{e}$ is a pure high extremal point.
We denote by $\{n_{p}^{(\ell)}\}_{\ell\in\mathbb{N}}$ the fixed point sequence, and let $\overline{m}_{\ell}$ be the number of pure high extremal points between $n_{p}^{(\ell)}$ and $n_{p}^{(\ell+1)}$ for all $\ell$. These notions not only help characterize the behavior of $J_{3}$, but also provide insight into its recursive and discrete piecewise linear structure, as illustrated in Figure 1.
Figure 1: Josephus function $J_{3}$ for $n\leq 50$.
Next, we present the main result of this paper, organized into three subsections. We begin by motivating the main goal of the manuscript by recalling the well-known explicit formula for the fixed points of $J_{2}$. Subsection 2.1 connects the search for fixed points of $J_{3}$ with the Chinese Remainder Theorem, representing each fixed point as a solution of a system of congruences with pairwise coprime moduli. Subsection 2.2 introduces a fractional number system in base $3/2$, following [1, 7], and studies its relevance to fixed points. Finally, Subsection 2.3 uses this system to recursively determine the digits in the base $3/2$ expansion of each fixed point of $J_{3}$.
2 Main Results and Motivations
In this paper, we aim to derive a recursive formula for the digits of consecutive fixed points for $J_{3}$ in a suitable base. Such a formula was obtained in full for the special case of $J_{2}$ in [6, 5]. In this case, the fixed points are given by the explicit formula $n_{p}^{(\ell)}=2^{\ell}-1$, for all $\ell\in\mathbb{N}$. Their expansions in base $2$ exhibit a particularly simple recursive pattern. Specifically, if
$n_{p}^{(\ell)}=(\hat{b}_{k}\cdots\hat{b}_{0})_{2},$
$\hat{b}_{i}\in\{0,1\}$ for all $i$, then the next fixed point satisfies
$n_{p}^{(\ell+1)}=(\hat{b}_{k}\cdots\hat{b}_{0}\,1)_{2},$
that is, the digit $1$ is appended to the right of the base $2$ expansion of $n_{p}^{(\ell)}$. Moreover, all digits $\hat{b}_{i}$ are ones, as shown in the following table:
$\ell$ $n_{p}^{(\ell)}=2^{\ell}-1$ Base $2$ Expansion
1 1 $(1)_{2}$
2 3 $(1\underline{\bf 1})_{2}$
3 7 $(11\underline{\bf 1})_{2}$
4 15 $(111\underline{\bf 1})_{2}$
5 31 $(1111\underline{\bf 1})_{2}$
6 63 $(11111\underline{\bf 1})_{2}$
7 127 $(111111\underline{\bf 1})_{2}$
8 255 $(1111111\underline{\bf 1})_{2}$
9 511 $(11111111\underline{\bf 1})_{2}$
10 1023 $(111111111\underline{\bf 1})_{2}$
Table 1: The first ten values of $n_{p}^{(\ell)}$ and their base $2$ expansions.
This elegant recurrence was the primary motivation for extending this result to $J_{3}$.
2.1 Fixed points of $J_{3}$ and the Chinese Remainder Theorem
We begin by recalling the recursive formula for the fixed point sequence $\{n_{p}^{(\ell)}\}_{\ell\in\mathbb{N}}$ of the function $J_{3}$, as stated in [2, Theorem 2]. Starting at $n_{p}^{(1)}=1$, we compute
$n_{p}^{(\ell+1)}$ $=\frac{3^{\overline{m}_{\ell}}(3n_{p}^{(\ell)}+2)-2^{\overline{m}_{\ell}}}{2^{\overline{m}_{\ell}+1}},$ (1a)
$\overline{m}_{\ell}$ $=\max\left\{m\in\mathbb{Z}_{+}\mid 2^{m}\text{ divides }3n_{p}^{(\ell)}+2\right\},$ (1b)
for every $\ell\in\mathbb{N}$. From equation (1a), we can derive, after a few algebraic manipulations, the equivalent identity
$2^{\overline{m}_{\ell}}(2n_{p}^{(\ell+1)}+1)=3^{\overline{m}_{\ell}}(3n_{p}^{(\ell)}+2),$ (2)
which tells us that $x_{1}=n_{p}^{(\ell+1)}$ and $x_{2}=n_{p}^{(\ell)}$ are each a solution to the following system of linear congruences:
$2x_{1}+1$ $\equiv 0\pmod{3^{\overline{m}_{\ell}}}$ (3a)
$3x_{2}+2$ $\equiv 0\pmod{2^{\overline{m}_{\ell}}}$. (3b)
Moreover, if we apply the same argument to $x_{1}=n_{p}^{(\ell+2)}$ and $x_{2}=n_{p}^{(\ell+1)}$, then they each satisfy:
$2x_{1}+1$ $\equiv 0\pmod{3^{\overline{m}_{\ell+1}}}$ (4a)
$3x_{2}+2$ $\equiv 0\pmod{2^{\overline{m}_{\ell+1}}}$. (4b)
Furthermore, since $n_{p}^{(\ell+1)}$ satisfies both (3a) and (4b), it must also satisfy the following system of congruences:
[…]
where $p:=\overline{m}_{\ell}$ and $q:=\overline{m}_{\ell+1}$. Note that when $pq=0$, our argument does not apply in this case. From this point on, we assume that $p,q\geq 1$. Independently, one can prove by induction on $p$ and $q$ that
[…]
Therefore, the fixed point $n_{p}^{(\ell+1)}$ satisfies the following system of linear congruences:
[…]
Since $3^{p}$ and $2^{q}$ are coprime positive integers, i.e., $\gcd(3^{p},2^{q})=1$, we can apply the classical form of the Chinese Remainder Theorem (CRT) (see [8, p. 294]). This guarantees a unique solution $z$ to (5) modulo $3^{p}\times 2^{q}$, given by
$z=a_{1}2^{q}y+a_{2}3^{p}x,$
where $x$ and $y$ are a pair of integers satisfying Bézout’s identity (see [8, p. 285]):
$3^{p}x+2^{q}y=\gcd(3^{p},2^{q})=1,$ (6)
where $\gcd$ is the greatest common divisor. In this way, we uncover an interesting connection between the sequence $\{n_{p}^{(\ell)}\}_{\ell\in\mathbb{N}}$ and the CRT. Moreover, since the explicit expression for $z$ depends on a solution to Bézout’s identity, it is natural to examine how $1$ can be expressed as a linear combination of $3^{p}$ and $2^{q}$.
$q$ Bézout’s identity
1 $1=3^{p}(1)+2^{1}\left(\tfrac{1}{2}(1-3^{p})\right),\quad\text{if }p\geq 1$
2 $1=\begin{cases}3^{p}(1)+2^{2}\left(\tfrac{1}{4}(1-3^{p})\right),&\text{if }p\equiv 0\pmod{2}\\[10.243pt] 3^{p}(-1)+2^{2}\left(\tfrac{1}{4}(1+3^{p})\right),&\text{if }p\equiv 1\pmod{2}\end{cases}$
3 $1=\begin{cases}3^{p}(1)+2^{3}\left(\tfrac{1}{8}(1-3^{p})\right),&\text{if }p\equiv 0\pmod{2}\\[10.243pt] 3^{p}(3)+2^{3}\left(\tfrac{1}{8}(1-3^{p+1})\right),&\text{if }p\equiv 1\pmod{2}\end{cases}$
4 $1=\begin{cases}3^{p}(1)+2^{4}\left(\tfrac{1}{16}(1-3^{p})\right),&\text{if }p\equiv 0\pmod{4}\\[10.243pt] 3^{p}(-5)+2^{4}\left(\tfrac{1}{16}(1+5\cdot 3^{p})\right),&\text{if }p\equiv 1\pmod{4}\\[10.243pt] 3^{p}(-7)+2^{4}\left(\tfrac{1}{16}(1+7\cdot 3^{p})\right),&\text{if }p\equiv 2\pmod{4}\\[10.243pt] 3^{p}(3)+2^{4}\left(\tfrac{1}{16}(1-3^{p+1})\right),&\text{if }p\equiv 3\pmod{4}\end{cases}$
[…]
Table 2: Bézout’s combinations for any $p\in\mathbb{N}$ and $q=1,2,3,4,5$.
Table 2 provides Bézout’s linear combinations for all $p\in\mathbb{N}$ and for $q=1,2,3,4,5$ in a compact form. We invite the interested reader to extend Table 2 to other values of $q$.
Remark.
To illustrate the preceding discussion, we present two representative examples from Table 3.
The first example corresponds to the following values $n_{p}^{(17)}=3986218$ and $n_{p}^{(18)}=102162424$ given in [3, Table 1] or Table 2, where $p=7$ (i.e., $\overline{m}_{17}$) and $q=1$ (i.e., $\overline{m}_{18}$), respectively. According to (5) and Table 2, $a_{1}=1093$, $a_{2}=0$, and $x=1$, y=1093y=-1093. Therefore,
$z=1093\times 2^{1}\times(-1093)+0\times 3^{7}\times 1=-2389298\equiv 3280\pmod{4374}$.
Thus, $n_{p}^{(18)}=102162424$ must be equivalent to 3280 modulo 4374, which is indeed true since (1021624243280)/4374=23356(102162424-3280)/4374=23356.
Another example is obtained when $n_{p}^{(13)}=46084$ and $n_{p}^{(14)}=103690$; see Table 3 below. In this case, $p=1$ (i.e., $\overline{m}_{13}=1$) and $q=5$ (i.e., $\overline{m}_{14}=5$). Here, $a_{1}=1$, $a_{2}=10$ in (5), and $x=11$, y=1y=-1 for the Bézout identity (6) since
$3^{1}\times 11+2^{5}\times(-1)=33-32=1.$
So,
$z=1\times 2^{5}\times(-1)+10\times 3^{1}\times 11=298\equiv 10\pmod{96}$.
Hence, $n_{p}^{(14)}=103690$ must be congruent to $10$ modulo $96$. Indeed, (10369010)/96=1080(103690-10)/96=1080.
2.2 Expansions in base $3/2$
We begin this subsection by noting that, by algebraic manipulations of the right-hand side of (1a), we obtain the following expression:
$n_{p}^{(\ell+1)}=\left(\frac{3}{2}\right)^{\overline{m}_{\ell}+1}n_{p}^{(\ell)}+\left(\frac{3}{2}\right)^{\overline{m}_{\ell}}-\frac{1}{2},$ (7)
which provides motivation for studying representations of natural numbers in base $3/2$. Inspired by (7) and the basic conceptual framework presented in the introduction to [7, Section 3], we move from base 3/2-3/2 to base $3/2$, which will allow us to characterize the fixed points of $J_{3}$ in a much simpler way. Here, we do not refer to the relatively well-known fractional number system in base $3/2$ with digits 0 or $1$. To recall how such a system works, let us choose the natural number $4$ and represent it in base $3/2$. Clearly, a good default approximation of this number is $1000.01001001_{3/2}$ because
$\left(\frac{3}{2}\right)^{3}+\left(\frac{3}{2}\right)^{-2}+\left(\frac{3}{2}\right)^{-5}+\left(\frac{3}{2}\right)^{-8}=3.99015012955.$
Instead of relying on that system, we aim to represent any positive integer $N$ in the form:
$N=\frac{1}{2}\left[d_{k}\left(\frac{3}{2}\right)^{k}+d_{k-1}\left(\frac{3}{2}\right)^{k-1}+\cdots+d_{1}\left(\frac{3}{2}\right)^{1}+d_{0}\right],$ (8)
where $d_{i}\in\mathcal{D}=\{0,1,2\}$ for some $k\in\mathbb{Z}_{+}$. If $k\geq 1$, we always assume that $d_{k}\neq 0$. When $N$ has the expansion of the form (8) we denote: $N=(d_{k}\cdots d_{0})_{3/2}.$ Note further that this expansion allows the digit $2$, which is not permitted in the standard base $3/2$ representation. Indeed, the expansion of the number $4$ in the form of (8) is $(212)_{3/2}$, since
$\frac{1}{2}\left[2\left(\frac{3}{2}\right)^{2}+1\left(\frac{3}{2}\right)^{1}+2\right]=4.$
We now describe the process for computing an expansion of the form (8). To this end, we use the following four-step inductive algorithm, presented in [1, p. 61]:
Step 1: Set $N_{0}=N$ and write $2N_{0}=3N_{1}+d_{0}$ with $d_{0}\in\mathcal{D}$ and $N_{1}\in\mathbb{Z}_{+}$. Note that $d_{0}$ is the unique digit in $\mathcal{D}$ satisfying the congruence $d_{0}={\rm mod}(2N_{0},3)$ and thus, $N_{1}$ is also uniquely determined.
Step 2: Define the non-negative integers $N_{1},N_{2},\ldots$ recursively by the equations $2N_{i}=3N_{i+1}+d_{i},$ with $d_{i}\in\mathcal{D}$, uniquely defined as $d_{i}={\rm mod}(2N_{i},3)$.
Step 3: We now prove by induction on $i$ that
$N=\left(\frac{3}{2}\right)^{i+1}N_{i+1}+\frac{1}{2}\left[d_{i}\left(\frac{3}{2}\right)^{i}+\cdots+d_{1}\left(\frac{3}{2}\right)+d_{0}\right].$ (9)
We verify the base case $i=0$: from Step 1,
$N=N_{0}=\frac{3}{2}N_{1}+\frac{1}{2}d_{0},$
which confirms that (9) holds for $i=0$. Now assume that (9) holds for $i=k\geq 0$. From Step 2, we have
$2N_{k+1}=3N_{k+2}+d_{k+1}.$
Substituting this into the inductive hypothesis gives us:
$N$ $=\left(\frac{3}{2}\right)^{k+1}N_{k+1}+\frac{1}{2}\left[d_{k}\left(\frac{3}{2}\right)^{k}+\cdots+d_{1}\left(\frac{3}{2}\right)+d_{0}\right]$
$=\left(\frac{3}{2}\right)^{k+1}\left(\frac{3}{2}N_{k+2}+\frac{1}{2}d_{k+1}\right)+\frac{1}{2}\left[d_{k}\left(\frac{3}{2}\right)^{k}+\cdots+d_{1}\left(\frac{3}{2}\right)+d_{0}\right]$
$=\left(\frac{3}{2}\right)^{k+2}N_{k+2}+\frac{1}{2}\left[d_{k+1}\left(\frac{3}{2}\right)^{k+1}+\cdots+d_{1}\left(\frac{3}{2}\right)+d_{0}\right],$
which proves that (9) holds for $i=k+1$. Thus, by induction, equation (9) holds for all $i\geq 0$.
Step 4: We show that $N_{\hat{i}}=0$ for some $\hat{i}$, ensuring that the algorithm terminates and yields a representation of the form (8).
From the recurrence relation
$N_{i+1}=\frac{2}{3}N_{i}-\frac{1}{3}d_{i}\leq\frac{2}{3}N_{i},$
the sequence $\{N_{i}\}$ is strictly decreasing whenever $N_{i}\geq 3$. This guarantees that, after finitely many steps, we will reach a value $N_{i_{0}}\in\{1,2\}$ for some $i_{0}\geq 1$. We consider the two possible cases: (i) If $N_{i_{0}}=2$ then $N_{i_{0}+1}=1$ and $N_{i_{0}+2}=0$, so the expansion terminates at $\hat{i}=i_{0}+2$. (ii) If $N_{i_{0}}=1$ then $N_{i_{0}+1}=0$, hence, $\hat{i}=i_{0}+1$.
Therefore, for any initial value $N=N_{0}\in\mathbb{N}$, the sequence $\{N_{i}\}$ eventually reaches zero, which confirms that $N$ has a finite expansion of the form (8). Since each digit $d_{i}$ is uniquely determined by the congruence $d_{i}\equiv 2N_{i}\pmod{3}$ and must lie in $\mathcal{D}=\{0,1,2\}$, it follows that the expansion is unique.
The next table shows the $3/2$ expansions of the type (8) for the first twenty fixed points of $J_{3}$.
$\ell$ $n_{p}^{(\ell)}$ $\overline{m}_{\ell}$ $3/2$ expansion
1 1 0 $(2)_{3/2}$
2 2 3 $(2\underline{\bf 1})_{3/2}$
3 13 0 $(21\underline{\bf 0112})_{3/2}$
4 20 1 $(210112\underline{\bf 1})_{3/2}$
5 46 2 $(2101121\underline{\bf 02})_{3/2}$
6 157 0 $(210112102\underline{\bf 012})_{3/2}$
7 236 1 $(210112102012\underline{\bf 1})_{3/2}$
8 532 1 $(2101121020121\underline{\bf 02})_{3/2}$
9 1198 2 $(210112102012102\underline{\bf 02})_{3/2}$
10 4045 0 $(21011210201210202\underline{\bf 012})_{3/2}$
11 6068 1 $(21011210201210202012\underline{\bf 1})_{3/2}$
12 13654 2 $(210112102012102020121\underline{\bf 02})_{3/2}$
13 46084 1 $(21011210201210202012102\underline{\bf 012})_{3/2}$
14 103690 5 $(21011210201210202012102012\underline{\bf 02})_{3/2}$
15 1181101 0 $(2101121020121020201210201202\underline{\bf 011112})_{3/2}$
16 1771652 1 $(2101121020121020201210201202011112\underline{\bf 1})_{3/2}$
17 3986218 7 $(21011210201210202012102012020111121\underline{\bf 02})_{3/2}$
18 102162424 1 $(2101121020121020201210201202011112102\underline{\bf 01111112})_{3/2}$
19 229865455 0 $(210112102012102020121020120201111210201111112\underline{\bf 02})_{3/2}$
20 344798183 0 $(21011210201210202012102012020111121020111111202\underline{\bf 1})_{3/2}$
Table 3: Values of $n_{p}^{(\ell)}$, $\overline{m}_{\ell}$, and their base $\frac{3}{2}$ expansions for $\ell=1,\ldots,20$.
2.3 Characterization of $\{n_{p}^{(\ell)}\}_{\ell\in\mathbb{N}}$ through the base $3/2$ number system
Observing the base $3/2$ expansions of the first twenty fixed points of $J_{3}$ (see Table 3), we are led to believe that there is a natural connection between these expansions and the divisibility condition that defines the sequence $\{\overline{m}_{\ell}\}_{\ell\in\mathbb{N}}$ (see equation (1b)). In particular, we have underlined some of the trailing digits of the expansions, as shown in Table 3.
Indeed, Table 3 exhibits several interesting features:
  1. 1.
    The base $3/2$ expansion of a fixed point $n_{p}^{(\ell)}$ repeats in the initial part of the base $3/2$ expansion of the next fixed point $n_{p}^{(\ell+1)}$.
  2. 2.
    The number of digits in the base $3/2$ expansion of $n_{p}^{(\ell+1)}$ is equal to the number of digits in the expansion of $n_{p}^{(\ell)}$ plus the number of pure high extremal points between them, $\overline{m}_{\ell}$, plus one.
  3. 3.
    The last $\overline{m}_{\ell}+1$ digits in the base $3/2$ expansion of $n_{p}^{(\ell+1)}$ are updated as follows:
    1. (a)
      The expansion is extended by the digit $1$ if $\overline{m}_{\ell}=0$.
    2. (b)
      The digits $02$ are appended to the expansion if $\overline{m}_{\ell}=1$.
    3. (c)
      The expansion is followed by the digit strings $012$, $0112$, $01112$, $\ldots$ if $\overline{m}_{\ell}=2,3,4,\ldots$.
We now state and prove the main result of this work, which confirms and formalizes these observations.
Theorem 1.
Assume that the base $3/2$ expansion of the fixed point $n_{p}^{(\ell)}$ of the Josephus function $J_{3}$ is given by $n_{p}^{(\ell)}=(\hat{d_{k}}\cdots\hat{d_{0}})_{3/2}.$ Then, the following statements hold:
If $\overline{m}_{\ell}=0$, then $n_{p}^{(\ell+1)}=(\hat{d_{k}}\cdots\hat{d_{0}}\,1)_{3/2}$.
If $\overline{m}_{\ell}=1$, then $n_{p}^{(\ell+1)}=(\hat{d_{k}}\cdots\hat{d_{0}}\,0\,2)_{3/2}$.
If $\overline{m}_{\ell}\geq 2$, then $n_{p}^{(\ell+1)}=(\hat{d_{k}}\cdots\hat{d_{0}}\,0\!\underbrace{1\cdots 1}_{\text{$(\overline{m}_{\ell}-1)$ ones}}\!2)_{3/2}$.
The total number of digits in the base $3/2$ expansion of the fixed point $n_{p}^{(\ell+1)}$ is equal to the number of digits in the base $3/2$ expansion of its predecessor, $n_{p}^{(\ell)}$, plus $\overline{m}_{\ell}+1$.
Proof.
To prove (i) let $\overline{m}_{\ell}=0$. Then, by (2),
$2n_{p}^{(\ell+1)}=3n_{p}^{(\ell)}+1.$
So, $d_{0}=1$, and in the induction formula (9), with $N=n_{p}^{(\ell+1)}$ and $N_{1}=n_{p}^{(\ell)}$ proving this case.
To prove (ii) and (iii), assume that $\overline{m}_{\ell}\geq 1$. We set $N_{0}=n_{p}^{(\ell+1)}.$ For simplicity, we define
$A_{i}=\dfrac{3n_{p}^{(\ell)}+2}{2^{\overline{m}_{\ell}-i}}$
for $i=0,1,\ldots,\overline{m}_{\ell}$. Note that the definition of $\overline{m}_{\ell}$ in (1b) implies that $A_{i}\in\mathbb{N}$ for all $i=0,1,\ldots,\overline{m}_{\ell}$. Then, by (2),
$2N_{0}=3(3^{\overline{m}_{\ell}-1}A_{0})-1=3(3^{\overline{m}_{\ell}-1}A_{0}-1)+2.$
So,
$d_{0}={\rm mod}(2N_{0},3)=2\quad\mbox{and}\quad N_{1}=3^{\overline{m}_{\ell}-1}A_{0}-1$
by observing the recursion (9). If $\overline{m}_{\ell}=1$,
$2N_{1}=2A_{0}-2=A_{1}-2=3n_{p}^{(\ell)}.$
Therefore, $d_{1}={\rm mod}(2N_{1},3)=0$. Hence, $d_{1}=0$ and the digits 2 and 0 are added to the right of the base $3/2$ expansion of $N_{2}=n_{p}^{(\ell)}$ to represent $n_{p}^{(\ell+1)}$, establishing item (ii).
If $\overline{m}_{\ell}\geq 2$, we can repeat the algorithm $\overline{m}_{\ell}-1$ times starting at:
Step 1: Here, we have
$2N_{1}=3(3^{\overline{m}_{\ell}-2}A_{1})-2=3(3^{\overline{m}_{\ell}-2}A_{1}-1)+1.$
Then,
$d_{1}={\rm mod}(2N_{1},3)=1\quad\mbox{and}\quad N_{2}=3^{\overline{m}_{\ell}-2}A_{1}-1.$
Until, Step $\overline{m}_{\ell}-1$: where we get,
$2N_{\overline{m}_{\ell}-1}=3A_{\overline{m}_{\ell}-1}-2=3(A_{\overline{m}_{\ell}-1}-1)+1.$
Hence,
$d_{\overline{m}_{\ell}-1}={\rm mod}(2N_{\overline{m}_{\ell}-1},3)=1\quad\mbox{and}\quad N_{\overline{m}_{\ell}}=A_{\overline{m}_{\ell}-1}-1.$
Finally, by applying the algorithm over again, we get
$2N_{\overline{m}_{\ell}}=2A_{\overline{m}_{\ell}-1}-2=A_{\overline{m}_{\ell}}-2=3n_{p}^{(\ell)}.$
Thus,
$d_{\overline{m}_{\ell}}=0\quad\mbox{and}\quad N_{\overline{m}_{\ell}+1}=n_{p}^{(\ell)}$
proving item (iii).
The item (iv) follows directly from (i), (ii), and (iii), as each case increases the digit count of the base $3/2$ expansion by exactly $\overline{m}_{\ell}+1$. ∎
This result is a natural generalization of the explicit formula for the fixed points of the Josephus function $J_{2}$, discussed at the beginning of this section. Moreover, [6, 5] show that the binary expansion of $n$ can also be used to compute $J_{2}(n)$ via a digit shift. Specifically, if $n=(b_{k}\cdots b_{0})_{2}$ and $j:=\max\left\{j\in\{k-1,\ldots,0\right\}\mid b_{j}\neq 0\}$, then $J_{2}(n)=(b_{j}\cdots b_{0}b_{k})_{2}$. It remains open whether: (i) the base $3/2$ expansions of $n$ and $n_{p}^{(\ell+1)}$, together with the information of $\overline{m}_{\ell}$, can be used to solve the Josephus problem, $J_{3}(n)$. Equations (9) and (10) in [2] may provide a method to determine the base $3/2$ expansion of $J_{3}(n)$; (ii) the fixed points of the Josephus function $J_{4}$ admit a recursive digit formula when expressed in base $4/3$ number system. Exploring such a representation could reveal structural patterns analogous to those observed in the base $3/2$ expansions of the fixed points of $J_{3}$, which in turn generalize the case of $J_{2}$ in base $2$.
3 Concluding Remarks
In this paper, we studied the fixed point sequence of the Josephus function $J_{3}$ and showed how its structure can be better understood through number theory and a non-integer base representation. First, we used the Chinese Remainder Theorem to describe each fixed point as the solution of a system of two linear congruences, exploiting the fact that the moduli involved are coprime. Then, we explored a numerical pattern that appears when the fixed points are expressed in base $3/2$ and provided a recursive formula for computing the digits of their expansions.
This connection offers a new and simple way to describe the fixed points of $J_{3}$ and opens the door to further analysis using tools from fractional number systems in this setting. Overall, the use of base $3/2$ provides a helpful perspective for understanding the behavior of the fixed points and suggests interesting directions for future research. One such direction is to investigate whether the fixed points of $J_{4}$ admit a recursive formula in base $4/3$, potentially revealing patterns similar to those found for $J_{3}$ and $J_{2}$. However, $J_{4}$ has two distinct pure high extremal points, which may require at least two indices to describe the number of pure high extremal points between consecutive fixed points.
Another open and promising direction is to determine whether the Josephus function itself, namely $J_{3}(n)$, can be computed directly using base $3/2$ representations. In this case, the expression for $J_{3}(n)$ given in (9) and (10) of [2] involves two indices, which complicates the derivation of a recursion for the base $3/2$ expansion of $J_{3}(n)$.
Acknowledgement. The first author would like to acknowledge the support provided by the National Science Foundation (NSF) grant # DMS - 2307328. The authors are grateful to the anonymous referees for their valuable comments and suggestions, which helped improve the quality of the paper.